### v1

@@ 3.1. Properties of the two-bridges partition
The relation between the lenght of interval is the classical dynamical partition and the two-bridges partition is the following: let $C>1$ and $\widehat{C}>1$ be the two constants given by real bounds of the classical and the two-bridges partition, respectively. If $I\in\mathcal{P}_{n}$ and $\widehat{I}\in\widehat{\mathcal{P}}_{n}$ and $I\cap\widehat{I}\neq\emptyset$ then
@@ 4. Idea of the proof
The proof of our Main Theorem consists in to construct a sequence of circle partitions that satisfy some especific properties. More precisely, we want to prove the following result:
@@
    the maximal lenght of an interval in $\mathcal{Q}_{n}^{f}$, and the maximal lenght of an interval in $\mathcal{Q}_{n}^{g}$, go to zero as $n$ goes to infinite;
@@
We will prove that the sequence of two-bridges partition $\{\mathcal{\widehat{P}}_{n}^{f}\}_{n\in\mathbb{N}}$, and its corresponding sequence $\{\mathcal{\widehat{P}}_{n}^{g}\}_{n\in\mathbb{N}}$ [^1] If the conjugacy $h$ identifies critical points of $f$ with critical points of $g$, then $h$ also identifies the corresponding two-bridges partitions $\mathcal{\widehat{P}}_{n}^{f}$ and $\mathcal{\widehat{P}}_{n}^{g}$, for all $n\in\mathbb{N}$. for $n\geq n_{0}$ where $n_{0}$ is given by Theorem 2.6, satisfy the three conditions in Proposition 4.1. The only condition that is not a consequence of the definition and properties of $\{\mathcal{\widehat{P}}_{n}^{f}\}_{n\in\mathbb{N}}$ is the third one, our main aim in this paper is to prove condition $(3)$ in Proposition 4.1.
@@
    First, we fixed $m\in\mathbb{N}$, and we the consider the two fundamental domain of the renormalizations $J_{m}(c_{0})$ and $J_{m}(c_{1})$. Inspired by [KT07], in Section 5 we prove that there exists an uniform strip around $m$ where, for all $n$ in that strip, the distance between each endpoint in $\widehat{E}_{n}$ (contained in $J_{m}(c_{0})$ or in $J_{m}(c_{1})$) and its image by $h$ is exponentially small, at rate $n$. We get those estimates in Proposition 6.1.
@@ 5. Un-bounded rotation number
From now on, for all $n\in\mathbb{N}$, we will renormalize around $c_{0}$, so we write $I_{n}^{f}$, $I_{n+1}^{f}$, $J_{n}^{f}$, $\mathfrak{c}_{n}$, $s_{n}^{f}$ and $\mathcal{R}^{n}f$ instead of $I_{n}^{f}(c_{0})$, $I_{n+1}^{f}(c_{0})$, $J_{n}^{f}(c_{0})$, $\mathfrak{c}_{n,0}^{f}$, $s_{n,0}^{f}$ and $\mathcal{R}_{0}^{n}f$ respectively. Since $h$ is the conjugacy that identifies corresponding critical points, then $\mathfrak{c}_{n}^{g}=h(\mathfrak{c}_{n})$. We notice that the construction, definitions and results given in this section can be extended if we renormalize around $c_{1}$.
We say that an irrational number $\rho=[a_{0},a_{1},\dots,a_{n},\dots]\in[0,1)$ is bounded type if there exists $A\in\mathbb{N}$ such that $\sup_{k}a_{k}\leq A$. The set of bounded type rotation numbers has zero Lebesgue measure in $[0,1]$. As we mentioned in the introduction, we focus our attention on un-bounded type rotation numbers, i.e. $\rho=[a_{0},a_{1},\cdots,a_{n},\cdots]$ such that there is no $A\in\mathbb{N}$ such that $\sup_{k}a_{k}<A$. In particular, the fundamental interval $I_{k}$ is subdivided by an un-bounded number of subintervals of the classical dynamical partition $\mathcal{P}_{k+1}$, and hence the graph of $\mathcal{R}^{k}f$ is arbitrarily close to the identity.
@@
Note that $S(\mathcal{R}^{n}f)<0$ in $[0,\mathcal{R}^{n}f(B_{n,f}(\mathfrak{c}_{n}))]\cup[(\mathcal{R}^{n}f)^{-1}(B_{n,f}(\mathfrak{c}_{n})),1]$. The proof follows from a small adaptation of the proof of [GMdM18, Lemma 6.1]. ∎
@@
Given $M>1$ there exists $L=L(M)\in\mathbb{N}$ such that $M_{n,L}^{f}$ is a non empty open interval, which can be an open interval or two open intervals, that contains at most two points $z_{n}^{f}$ and $w_{n}^{f}$ which attains the minimum of $|z-\mathcal{R}^{n}f(z)|$ inside $M_{n,L}^{f}$. For those points we have $D\mathcal{R}^{n}f(z_{n}^{f})=1$, $D\mathcal{R}^{n}f(w_{n}^{f})=1$, $D^{2}\mathcal{R}^{n}f(z_{n}^{f})<-1/M$ and $D^{2}\mathcal{R}^{n}f(w_{n}^{f})<-1/M$.
If the points $z_{n}^{f}$ and $w_{n}^{f}$ in Corollary 5.2 exist, then they are called the centers of $M_{n,L}^{f}$. The exponential convergence of the renormalizations, in the $C^{1}$-topology, implies that
@@ 5.1. Tubular coordinates
We are interested in obtain asymptotics estimates of $\mathcal{F}_{m}$ and of $\mathcal{G}_{m}$, near the origin. Since $\mathcal{F}_{m}^{\prime}(0)=1$, $\mathcal{F}_{m}^{\prime\prime}(0)=2$ and $\varepsilon_{m}=\mathcal{F}_{m}(0)=\min_{x}\{\mathcal{F}_{m}(x)-x\}<0$, then there exists $\mathfrak{C}>0$ such that
@@
Let us give some more know estimates. First, let $i_{c},\widehat{i_{c}}\in\mathbb{N}$ the smallest natural number such that $\mathcal{F}_{m}^{i_{c}}(\phi_{m}(1))\in(0,-\varepsilon_{m}]$ and $\mathcal{F}_{m}^{-\widehat{i_{c}}}(\phi_{m}(B_{m,f}(\mathfrak{c}_{m})))\in(0,-\varepsilon_{m}]$, respectively. Moreover, let $i_{r},i_{\ell}\in\mathbb{N}$ be such that $\mathcal{F}_{m}^{i_{r}}(\phi_{m}(1))$ is the most right endpoint of the Funnel and $\mathcal{F}_{m}^{i_{\ell}}(\phi_{m}(1))$ is the most left endpoint of the Funnel. Similarly, let $\widehat{i_{r}},\widehat{i_{\ell}}\in\mathbb{N}$ such that $\mathcal{F}_{m}^{-\widehat{i_{\ell}}}(\phi_{m}(B_{m,f}(\mathfrak{c}_{m})))$ is the most left endpoint of the Funnel and $\mathcal{F}_{m}^{-\widehat{i_{r}}}(\phi_{m}(B_{m,f}(\mathfrak{c}_{m})))$ is the most right endpoint of the Funnel.
@@
Note that $i_{\kappa}\neq j_{\kappa}$ for $\kappa\in\{c,r,\ell\}$. Next result follows from [KT07, Pages 206-207].
@@
There exists constants $D_{5}$ and $D_{6}$ such that if $r(m)\geq D_{6}$ then
@@
We also have a similar result when $\ell(m)$ is bigger enough.
@@ 6. Key estimates for points in $\widehat{E}_{n+1}$
$\leq\left|\mathcal{R}^{m}f(0)(B_{m+1,g}(h(v))-B_{m+1,f}(v))\right|+|B_{m+1,f}(v)|\left|\mathcal{R}^{m}g(0)-\mathcal{R}^{m}f(0)\right|$
@@
If $z_{m}\notin B_{m,f}([f^{-(L+1)q_{m+1}}(\mathfrak{c}_{m}),f^{-(L+1+N)q_{m+1}}(\mathfrak{c}_{m})])\cap B_{m,f}([f^{q_{m}+(L+P)q_{m+1}}(c_{0}),f^{q_{m}+Lq_{m+1}}(c_{0})])$, then since $D\mathcal{R}^{m}f(x)<1$ in $B_{m,f}([f^{q_{m}+(L+P)q_{m+1}}(c_{0}),f^{q_{m}+Lq_{m+1}}(c_{0})])$ we have, for $1\leq j\leq P$ and $\lambda_{1}=\max\{\lambda,\lambda^{1/d}\}$
@@
and since $D\mathcal{R}^{m}f(x)>1$ in $B_{m,f}([f^{-(L+1)q_{m+1}}(\mathfrak{c}_{m}),f^{-(L+1+N)q_{m+1}}(\mathfrak{c}_{m})])$ then, for $1\leq j\leq P$, we have
@@
$0<|\varepsilon_{m}|=|\mathcal{F}(0)|<\frac{|D^{2}\mathcal{R}^{m}f(z_{m})|}{2\,r(m)}<\frac{M}{4(L+P)}<\frac{M}{4P}<\frac{M}{4}\,\lambda^{m/2}.$ (10)
@@
$|\phi_{m}(B_{m,f}(v))|\leq C_{0}\,|\varepsilon_{m}|^{1/2+\alpha}\leq C_{0}\,\left(\frac{M}{4}\right)^{1/2+\alpha}\,\left(\lambda^{1/2+\alpha}\right)^{m/2},$
@@
$|B_{m,f}(v)-z_{m}|\leq\frac{2C_{0}}{M}\left(\frac{M}{4}\right)^{1/2+\alpha}\,\left(\lambda^{1/2+\alpha}\right)^{m/2}.$
Otherwise, if $\phi_{m}(B_{m,f}(v))$ is inside the Funnel, then $B_{m,f}(v)$ corresponds to some iterate $\mathcal{F}_{m}^{i}(\phi(1))$ or $\mathcal{F}^{-i}(\phi(B_{m,f}(\mathfrak{c}_{m})))$ for some $i$. In the first case, using inequalities (10) and (5) for the map $\mathcal{F}$, with $s_{0}=\mathcal{F}^{i_{c}}(\phi(1))$ and $s_{k}=\mathcal{F}^{i_{c}-k}(\phi(1))$, we get, for all $k\geq 0$
@@
and we get the claim for any endpoint $v\in\widehat{E}_{m+1}$ such that $\phi(B_{m,f}(v))$ is inside the left part of the Funnel. Moreover, since any endpoint $v\in\widehat{E}_{m+1}$ which is inside the right part of the Funnel is between consecutive points of the form $f^{kq_{m+1}+q_{m}}(c_{0})$ or $f^{-kq_{m+1}}(\mathfrak{c}_{m})$, then we get the claim also for all these points. ∎
@@
Let $m=n+1$. Note that the set $\widehat{E}_{n+1}\cap J_{n+1}$ is equal to $\{f^{q_{n+1}}(c_{0}),f^{q_{n+2}}(c_{0}),\mathfrak{c}_{n}^{f}\}$ or $\{f^{q_{n+1}}(c_{0}),f^{q_{n+2}}(c_{0})\}$ depending on whether $m$ is or is not a two-bridges level. We observe that $|B_{n+1,f}(f^{q_{n+1}}(c_{0}))-B_{n+1,g}(g^{q_{n+1}}(c_{0}))|=0$, that inequality (9) implies $|B_{n+1,f}(\mathfrak{c}_{n}^{f})-B_{n+1,g}(\mathfrak{c}_{n}^{g})|\leq C\,\lambda^{(n+1)/d}$ and that inequality (LABEL:eq:carlesonpaso1) implies that $|B_{n+1,f}(f^{q_{n+2}}(c_{0}))-B_{n+1,g}(g^{q_{n+2}}(c_{0}))|\leq C\,\lambda^{n+1}$. Therefore,
@@
If $m<n+1$ then by previous inequality, Lemma 6.1 (inductively) and the fact that $(1-b_{1})n\leq m$, we get for $\widetilde{C}=\max\{C,C_{1}\}>1$

### v2

@@ 3.1. Properties of the two-bridges partition
The relation between the length of interval is the classical dynamical partition and the two-bridges partition is the following: let $C>1$ and $\widehat{C}>1$ be the two constants given by real bounds of the classical and the two-bridges partition, respectively. If $I\in\mathcal{P}_{n}$ and $\widehat{I}\in\widehat{\mathcal{P}}_{n}$ and $I\cap\widehat{I}\neq\emptyset$ then
@@ 4. Idea of the proof
The proof of our Main Theorem consists in to construct a sequence of circle partitions that satisfy some specific properties. More precisely, we want to prove the following result:
@@
    the maximal length of an interval in $\mathcal{Q}_{n}^{f}$, and the maximal length of an interval in $\mathcal{Q}_{n}^{g}$, go to zero as $n$ goes to infinite;
@@
We will prove that the sequence of two-bridges partition $\{\mathcal{\widehat{P}}_{n}^{f}\}_{n\in\mathbb{N}}$, and its corresponding sequence $\{\mathcal{\widehat{P}}_{n}^{g}\}_{n\in\mathbb{N}}$ [^1] If the conjugacy $h$ identifies critical points of $f$ with critical points of $g$, then $h$ also identifies the corresponding two-bridges partitions $\mathcal{\widehat{P}}_{n}^{f}$ and $\mathcal{\widehat{P}}_{n}^{g}$, for all $n\in\mathbb{N}$. for $n\geq n_{0}$ where $n_{0}$ is given by Theorem 2.6, satisfy the three conditions of Proposition 4.1. The only condition that is not a immediate consequence of the definition and properties of $\{\mathcal{\widehat{P}}_{n}^{f}\}_{n\in\mathbb{N}}$ is the third one, the main technical result in this paper is that condition $(3)$ of Proposition 4.1 holds.
@@
    First, we fixed $m\in\mathbb{N}$, and we the consider the two fundamental domain of the renormalizations $J_{m}(c_{0})$ and $J_{m}(c_{1})$. Inspired by [KT07], in Section 5 we prove that there exists an uniform strip around $m$ where, for all $n$ in that strip, the distance between each endpoint in $\widehat{E}_{n}$ (contained in $J_{m}(c_{0})$ or in $J_{m}(c_{1})$) and its image by $h$ is exponentially small, with rate $n$. We get those estimates in Proposition 6.1.
@@ 5. Un-bounded rotation number
From now on, for all $n\in\mathbb{N}$, we will renormalize around $c_{0}$, so we write $I_{n}^{f}$, $I_{n+1}^{f}$, $J_{n}^{f}$, $\mathfrak{c}_{n}$, $s_{n}^{f}$ and $\mathcal{R}^{n}f$ instead of $I_{n}^{f}(c_{0})$, $I_{n+1}^{f}(c_{0})$, $J_{n}^{f}(c_{0})$, $\mathfrak{c}_{n,0}^{f}$, $s_{n,0}^{f}$ and $\mathcal{R}_{0}^{n}f$ respectively. Since $h$ is the conjugacy that identifies corresponding critical points, then $\mathfrak{c}_{n}^{g}=h(\mathfrak{c}_{n})$. We notice that the construction, definitions and results given in this section also apply if we renormalize around $c_{1}$.
We say that an irrational number $\rho=[a_{0},a_{1},\dots,a_{n},\dots]\in[0,1)$ is bounded type if there exists $A\in\mathbb{N}$ such that $\sup_{k}a_{k}\leq A$. The set of bounded type rotation numbers has zero Lebesgue measure in $[0,1]$. As we mentioned in the introduction, we focus our attention on un-bounded type rotation numbers, i.e. $\rho=[a_{0},a_{1},\cdots,a_{n},\cdots]$ such that there is no $A\in\mathbb{N}$ such that $\sup_{k}a_{k}<A$. In particular, the fundamental interval $I_{k}$ is subdivided by an un-bounded number of sub-intervals of the classical dynamical partition $\mathcal{P}_{k+1}$, and hence the graph of $\mathcal{R}^{k}f$ is arbitrarily close to the identity.
@@
Note that $S(\mathcal{R}^{n}f)<0$ in $[0,\mathcal{R}^{n}f(B_{n,f}(\mathfrak{c}_{n}))]\cup[(\mathcal{R}^{n}f)^{-1}(B_{n,f}(\mathfrak{c}_{n})),1]$. The proof then follows from a small adaptation of the proof of [GMdM18, Lemma 6.1]. ∎
@@
Given $M>1$ there exists $L=L(M)\in\mathbb{N}$ such that $M_{n,L}^{f}$ is a non empty set, which can be an open interval or two open intervals, that contains at most two points $z_{n}^{f}$ and $w_{n}^{f}$ which attains the minimum of $|z-\mathcal{R}^{n}f(z)|$ inside $M_{n,L}^{f}$. For those points we have $D\mathcal{R}^{n}f(z_{n}^{f})=1$, $D\mathcal{R}^{n}f(w_{n}^{f})=1$, $D^{2}\mathcal{R}^{n}f(z_{n}^{f})<-1/M$ and $D^{2}\mathcal{R}^{n}f(w_{n}^{f})<-1/M$.
If the points $z_{n}^{f}$ and $w_{n}^{f}$ in Corollary 5.2 exist, then they are called the centers of $M_{n,L}^{f}$. The exponential convergence of the renormalizations, in the $C^{2}$-topology, implies that
@@ 5.1. Tubular coordinates
We are interested in obtaining asymptotic estimates of $\mathcal{F}_{m}$ and of $\mathcal{G}_{m}$, near the origin. Since $\mathcal{F}_{m}^{\prime}(0)=1$, $\mathcal{F}_{m}^{\prime\prime}(0)=2$ and $\varepsilon_{m}=\mathcal{F}_{m}(0)=\min_{x}\{\mathcal{F}_{m}(x)-x\}<0$, then there exists $\mathfrak{C}>0$ such that
@@
Let us give some more known estimates. First, let $i_{c},\widehat{i_{c}}\in\mathbb{N}$ the smallest natural number such that $\mathcal{F}_{m}^{i_{c}}(\phi_{m}(1))\in(0,-\varepsilon_{m}]$ and $\mathcal{F}_{m}^{-\widehat{i_{c}}}(\phi_{m}(B_{m,f}(\mathfrak{c}_{m})))\in(0,-\varepsilon_{m}]$, respectively. Moreover, let $i_{r},i_{\ell}\in\mathbb{N}$ be such that $\mathcal{F}_{m}^{i_{r}}(\phi_{m}(1))$ is the most right endpoint of the Funnel and $\mathcal{F}_{m}^{i_{\ell}}(\phi_{m}(1))$ is the most left endpoint of the Funnel. Similarly, let $\widehat{i_{r}},\widehat{i_{\ell}}\in\mathbb{N}$ such that $\mathcal{F}_{m}^{-\widehat{i_{\ell}}}(\phi_{m}(B_{m,f}(\mathfrak{c}_{m})))$ is the most left endpoint of the Funnel and $\mathcal{F}_{m}^{-\widehat{i_{r}}}(\phi_{m}(B_{m,f}(\mathfrak{c}_{m})))$ is the most right endpoint of the Funnel.
@@
Note that $i_{\kappa}\neq j_{\kappa}$ for $\kappa\in\{c,r,\ell\}$. Next result follows from a small adaptation of [KT07, Pages 206-207].
@@
There exists a constant $D_{5}>0$ such that if $r(m)$ is sufficiently big then
@@
We also have a similar result when $\ell(m)$ is sufficiently big.
@@ 6. Key estimates for points in $\widehat{E}_{n+1}$
$\leq\left|\mathcal{R}^{m}f(0)(B_{m+1,g}(h(v))-B_{m+1,f}(v))\right|+|B_{m+1,g}(h(v))|\left|\mathcal{R}^{m}g(0)-\mathcal{R}^{m}f(0)\right|$
@@
If $z_{m}\notin B_{m,f}([f^{-(L+1)q_{m+1}}(\mathfrak{c}_{m}),f^{-(L+1+P)q_{m+1}}(\mathfrak{c}_{m})])\cap B_{m,f}([f^{q_{m}+(L+P)q_{m+1}}(c_{0}),f^{q_{m}+Lq_{m+1}}(c_{0})])$, then since $D\mathcal{R}^{m}f(x)<1$ in $B_{m,f}([f^{q_{m}+(L+P)q_{m+1}}(c_{0}),f^{q_{m}+Lq_{m+1}}(c_{0})])$ we have, for $1\leq j\leq P$ and $\lambda_{1}=\max\{\lambda,\lambda^{1/d}\}$
@@
and since $D\mathcal{R}^{m}f(x)>1$ in $B_{m,f}([f^{-(L+1)q_{m+1}}(\mathfrak{c}_{m}),f^{-(L+1+P)q_{m+1}}(\mathfrak{c}_{m})])$ then, for $1\leq j\leq P$, we have
@@
$0<|\varepsilon_{m}|=|\mathcal{F}_{m}(0)|<\frac{|D^{2}\mathcal{R}^{m}f(z_{m})|}{2\,r(m)}<\frac{M}{4(L+P)}<\frac{M}{4P}<\frac{M}{4}\,\lambda^{m/2}.$ (11)
@@
$|\phi_{m}(B_{m,f}(v))|\leq C_{0}\,|\varepsilon_{m}|^{\frac{1}{2+\alpha}}\leq C_{0}\,\left(\frac{M}{4}\right)^{\frac{1}{2+\alpha}}\,\left(\lambda^{\frac{1}{2+\alpha}}\right)^{m/2},$
@@
$|B_{m,f}(v)-z_{m}|\leq\frac{2C_{0}}{M}\left(\frac{M}{4}\right)^{\frac{1}{2+\alpha}}\,\left(\lambda^{\frac{1}{2+\alpha}}\right)^{m/2}.$
Otherwise, if $\phi_{m}(B_{m,f}(v))$ is inside the Funnel, then $B_{m,f}(v)$ corresponds to some iterate $\mathcal{F}_{m}^{i}(\phi(1))$ or $\mathcal{F}_{m}^{-i}(\phi(B_{m,f}(\mathfrak{c}_{m})))$ for some $i$. In the first case, using inequalities (11) and (6) for the map $\mathcal{F}_{m}$, with $s_{0}=\mathcal{F}_{m}^{i_{c}}(\phi(1))$ and $s_{k}=\mathcal{F}_{m}^{i_{c}+k}(\phi(1))$, we get, for all $k\geq 0$
@@
and we get the claim for any endpoint $v\in\widehat{E}_{m+1}$ such that $\phi_{m}(B_{m,f}(v))$ is inside the left part of the Funnel. Moreover, since any endpoint $v\in\widehat{E}_{m+1}$ which is inside the right part of the Funnel is between consecutive points of the form $f^{kq_{m+1}+q_{m}}(c_{0})$ or $f^{-kq_{m+1}}(\mathfrak{c}_{m})$, then we get the claim also for all these points. ∎
@@
Let $m=n+1$. Note that the set $\widehat{E}_{n+1}\cap J_{n+1}$ is equal to $\{f^{q_{n+1}}(c_{0}),f^{q_{n+2}}(c_{0}),\mathfrak{c}_{n}^{f}\}$ or $\{f^{q_{n+1}}(c_{0}),f^{q_{n+2}}(c_{0})\}$ depending on whether $m$ is or is not a two-bridges level. We observe that (by definition) $|B_{n+1,f}(f^{q_{n+1}}(c_{0}))-B_{n+1,g}(g^{q_{n+1}}(c_{0}))|=0$, Inequality (10) implies that $|B_{n+1,f}(\mathfrak{c}_{n}^{f})-B_{n+1,g}(\mathfrak{c}_{n}^{g})|\leq C\,\lambda^{(n+1)/d}$ and that the hypotheses under the renormalizations (Inequality (1)) implies that $|B_{n+1,f}(f^{q_{n+2}}(c_{0}))-B_{n+1,g}(g^{q_{n+2}}(c_{0}))|\leq C\,\lambda^{n+1}$. Therefore,
@@
If $m<n+1$ then by previous inequality and Lemma 6.1 (inductively) we get for $\widetilde{C}=\max\{C,C_{1}\}>1$